# Systems with Switching Causal Relations: A Meta-Causal Perspective

### v1

@@ 1 Introduction
Structural Causal Models (SCM) have become the de facto formalism in causality by representing causal relations a structural equation models (Pearl (2009); Spirtes et al. (2000)). While SCM are most useful for representing the intricate mechanistic details of systems, it can be challenging to derive the general qualitative behavior that emerges from the interplay of individual equations. Talking about the general type of relation emitted by particular mechanisms generalizes above the narrow computational view and has the ability to inspect causal systems from a more general perspective. Causal graphs can be subject to change whenever novel mechanisms emerge or vanish within a system. Prominently, agents can ‘break’ the natural unfolding of systems dynamics by forecasting system behavior and preemptively intervening in the course of events. As inherent parts of the environment, agents commonly establish or suppress the emergence of causal connections (Zhang & Bareinboim, 2017; Lee & Bareinboim, 2018; Dasgupta et al., 2019).
Consider the scenario shown in Fig. 1 (left), where an agent $A$ (with position $A_{X}$) follows an agent $B$ (with position $B_{X}$) according to its policy $A_{\operatorname{\bm{\pi}}}$. We are interested in answering the question ‘What is the cause of the current position of agent A?’. In general, the observed system can be formalized as follows: $A_{X}:=f_{A}(B_{X},U_{A})$ and $B_{X}:=f_{B}(U_{B})$. From a classical causal perspective we observe that $B\rightarrow A$, since $A_{X}$ is conditioned on $B_{X}$ via the equation $f_{A}$. As a result, we infer that $B_{X}$ is the cause of $A$’s position. This perspective, only attributes causes within the current state of the system. However, it disregards the causes that lead to the emergence of $f_{A}$. Only when we take a meta-causal stance and think about how the equation $f_{A}$ came to be in the first place, we can give the answer that $A$, through its policy $A_{\operatorname{\bm{\pi}}}$, is the (meta-causal) cause for its own position. [^1] Although similar conclusions might be derived from classical causal perspectives, we refer the reader to our discussions in Sec. 4 for further considerations and limitations on these approaches.
@@ 2 Background
Causal Models. A common formalization of causality is provided via Structural Causal Models (SCM; (Spirtes et al., 2000; Pearl, 2009)). An SCM is defined as a tuple $\operatorname{\mathcal{M}}:=(\operatorname{\mathbf{U}},\operatorname{\mathbf{V}},\operatorname{\mathbf{F}},P_{\operatorname{\mathbf{U}}})$, where $\operatorname{\mathbf{U}}$ is the set of exogenous variables, $\operatorname{\mathbf{V}}$ is the set of endogenous variables, $\operatorname{\mathbf{F}}$ is the set of structural equations determining endogenous variables, and $P_{\operatorname{\mathbf{U}}}$ is the distribution over exogenous variables $\operatorname{\mathbf{U}}$. Every endogenous variable $V_{i}\in\operatorname{\mathbf{V}}$ is determined via a structural equation $v_{i}:=f_{i}(\text{pa}(v_{i}))$ that takes in a set of parent values $\text{pa}(v_{i})$, consisting of endogenous and exogenous variables, and outputs the value of $v_{i}$. The set of all variables is denoted by $\operatorname{\mathbf{X}}=\operatorname{\mathbf{U}}\cup\operatorname{\mathbf{V}}$ with values $\mathbf{x}\in\operatorname{\bm{\mathcal{X}}}$ and $N=|\operatorname{\mathbf{X}}|$. Every SCM $\operatorname{\mathcal{M}}$ entails a directed graph $\mathcal{G}^{\operatorname{\mathcal{M}}}$ that can be constructed by adding edges $(i,j)\in\operatorname{\mathbf{X}}\times\operatorname{\mathbf{X}}$ for all variables $X_{j}\in\operatorname{\mathbf{X}}$ and its parents $X_{i}\in\text{Pa}(X_{j})$. This can be expressed as an adjacency matrix $A\in\mathbb{B}^{N\times N}$ where $A_{ij}:=1$ if $(i,j)\in\mathcal{G}$ and $A_{ij}:=0$ otherwise.
@@
Mediation Processes. Causal effects are embedded in an environment that governs the dynamics and mediates between different causes and effects. To reason about the existence of causal relations, we need to know and consider this process-embedding environment. We define a mediation process $\operatorname{\mathcal{E}}=(\mathcal{S},\operatorname{\sigma})$, adapted from Markov Decision Processes (MDP; Bellman (1957)) for our setting. Here, $\mathcal{S}$ is the state space of the environment and $\operatorname{\sigma}:\mathcal{S}\rightarrow\mathcal{S}$ is the (possibly non-deterministic) transition function that takes the current state and outputs the next one. If we also have an initial state $\operatorname{s_{0}}\in\mathcal{S}$, we call $(\operatorname{\mathcal{E}},\operatorname{s_{0}})$ an initialized mediation process. As we are concerned with the general mediation process, we omit the common notion of a reward function $r$. Furthermore, we omit an explicit action space $\mathcal{A}$ and agents’ policy $\operatorname{\bm{\pi}}$ and model actions directly as part of the transition function $\operatorname{\sigma}$. In accordance to considerations of (Cohen, 2022), this eases our treatise of environment processes and agent actions as now both are defined over the same domain.
@@
Typing. In the following section, we will make use of an identification function $\operatorname{\mathcal{I}}$ to determine the presence or absence of edges between any two variables. In particular, one could make use of a set of different identification functions to identify different types of edges. Previous work on typing causality exists (Brouillard et al., 2022), but is primarily concerned with types of variables, rather than the typing of structural relations. Several other works in the field of cognitive science consider the perception of different types of mechanistic relations (e.g., ‘causing’, ‘enabling’, ‘preventing’, ‘despite’) based on the role that different objects play in physical scenes (Chockler & Halpern, 2004; Wolff, 2007; Sloman et al., 2009; Walsh & Sloman, 2011; Gerstenberg, 2022). Since all objects are usually governed by the same physical equations, this assignment of types serves to provide post hoc explanations of a scene, rather than to identify inherent properties of its computational aspects. Gerstenberg (2024) provides an ‘intuitive psychology’ example of agents that fits well with our scenario.
@@ 3 Meta-Causality
Meta-Causality is concerned with the general mechanisms that lead to the emergence of causal relations. Usually, the mediating process may be too fine-grained to yield interpretable models. Therefore, we consider a set of variables of interest $\operatorname{\mathbf{X}}$ derived via $\varphi:\operatorname{\mathcal{S}}\to\operatorname{\bm{\mathcal{X}}}$. Here, $\varphi$ could be defined as a summarization or abstraction operation over the state space (Rubenstein et al., 2017; Beckers & Halpern, 2019; Anand et al., 2022; Wahl et al., 2023; Kekić et al., 2023; Willig et al., 2023). In order to identify the type of causal relations from a mediating process, we need to be able to decide on what constitutes such a type.
@@
For a given mediation process $\operatorname{\mathcal{E}}=(\operatorname{\mathcal{S}},\operatorname{\sigma})$ a meta-causal frame is a tuple $\operatorname{\mathcal{F}}=(\operatorname{\mathcal{E}},\operatorname{\mathbf{X}},\operatorname{\tau},\operatorname{\mathcal{I}})$ with:
  • a type-encoder $\operatorname{\tau}:\operatorname{\mathcal{X}}_{i}\times\operatorname{\mathcal{X}}^{\operatorname{\mathcal{S}}}_{j}\to\operatorname{\mathcal{T}}$ that determines the relationship between the value of a causal variable $\operatorname{\mathcal{X}}_{i}$ and the change of another causal variable $\operatorname{\mathcal{X}}^{\operatorname{\mathcal{S}}}_{j}=\{\psi:\operatorname{\mathcal{S}}\to\operatorname{\mathcal{X}}_{j}\}$ with respect to a environmental state space $\operatorname{\mathcal{S}}$ which we call a type $t\in\operatorname{\mathcal{T}}$,
  • an identification function $\operatorname{\mathcal{I}}:\operatorname{\mathcal{S}}\times\operatorname{\mathbf{X}}\times\operatorname{\mathbf{X}}\rightarrow\operatorname{\mathcal{T}}$ with $\operatorname{\mathcal{I}}(s,X_{i},X_{j})\mapsto t:=\operatorname{\tau}(\varphi(s)_{i},\varphi_{j}\circ\operatorname{\sigma})$ that takes a state of the environment and assigns a type for every pair of causal variables $X_{i},X_{j}\in\operatorname{\mathbf{X}}$.
Types generalize the role of edges in causal graphs. In most classical scenarios, the co-domain $\operatorname{\mathcal{T}}$ of the type encoder $\operatorname{\tau}$ is chosen to be Boolean, representing the existence or absence of edges. In other cases, different values $t\in\mathcal{T}$ can be understood as particular types of edges, like positive, negative, or the absence of influence. This will help us to distinguish meta-causal states that share the same graph adjacency. The only requirement for $\operatorname{\mathcal{T}}$ is that it must contain a special value $\operatorname{\mathit{0}}$, which indicates the total absence of an edge. Meta-causal states now generalize the idea of binary adjacency matrices.
@@
In a meta-causal frame $\operatorname{\mathcal{F}}=(\operatorname{\mathcal{E}},\operatorname{\mathbf{X}},\operatorname{\tau},\operatorname{\mathcal{I}})$, a meta-causal state is a matrix $T\in\operatorname{\mathcal{T}}^{N\times N}$. For a given environment state $s\in\operatorname{\mathcal{S}}$, the actual meta-causal state $T_{s}$ has the entries $T_{s,ij}:=\operatorname{\mathcal{I}}(s,X_{i},X_{j})=\operatorname{\tau}(\varphi(s)_{i},\varphi_{j}\circ\operatorname{\sigma})$.
@@
Usually, we have the objective to learn the transition function $\delta$ for an unknown state transition function $\operatorname{\sigma}$. The state transition function $\delta$ can be approximated as $\delta(T,s):=\operatorname{\mathcal{I}}(\operatorname{\sigma}^{\prime}(T,s),X_{i},X_{j})=\operatorname{\tau}(\varphi(s)_{i},\varphi_{j}\circ\operatorname{\sigma}^{\prime}(T,s))$.
@@
Inferring Meta-Causal States. Even if the state transition function is known, it may be unclear from a single observation which exact meta-causal state led to the generation of a particular observation $S$. This is especially the case when two different meta-causal states can fit similar environmental dynamics. Even in the presence of latent factors (e.g., an agent’s internal policy), the current dynamics of a system (e.g., induced by the agent’s current policy) can sometimes be inferred from a series of observed environment states. This requires knowledge of the meta-causal dynamics, and is subject to the condition that sequences of observed states uniquely characterize the meta-causal state. Since MCM are defined as finite state machines, the exact condition for identifying the meta-causal state is that observed sequences are homing sequences (Moore et al., 1956). Note that the following example of a game of tag presented below exactly satisfies this condition, where the meta-causal states produce disjunct sets of environment states (‘agent A faces agent B’, or ‘agent B faces agent A’), and thus can be inferred from either a single observation, when movement directions are observed, or two observations, when they need to be inferred from the change in position of two successive observations.
@@
$(t_{A\rightarrow B}=\text{chasing})\iff(\dot{A}_{pos}\bm{\cdot}(B_{pos}-A_{pos})>0)$
@@
Causal and Anti-Causal Considerations of Meta-Causal States. Assigning meta-causal states to particular system observations can be understood as labeling the individual observations. However, it is generally unclear whether the meta-causal state has an observational or a controlling character on the system under consideration. One could ask the question whether the system dynamics cause the meta-causal state, whether the meta-causal state causes the system dynamics, (or whether they are actually the same concept), similar to the well-known discussion “On Causal and Anticausal Learning” by Schölkopf et al. (2012), but from a meta-causal perspective.
@@
Role of Contextual Independencies. Changes in the causal graph are often times attributed to changes in the environment, as for example prominently leveraged in invariant causal prediction of Peters et al. (2016); Heinze-Deml et al. (2018). In general, the effects of a changing causal graph can be attributed to the change in some external factor acting on the system. Under certain circumstances, we can indeed represent meta-causal models via standard SCM conditioned on some external factor $Z\in\mathcal{Z}$. Suppose there is a surjective function $\psi:\mathcal{Z}\to\operatorname{\mathcal{T}}^{N\times N}$ that can map from the $Z$ to every meta causal state $\operatorname{T}$. As a result, one can define a new set of structural equations $f^{\prime}_{i}:=\bigcirc_{i\in 1..N}f^{\prime}_{ij}$ and
$f_{ij}|Z:=\begin{cases}f^{\psi(z)}_{ij}&\textbf{if~{}}\psi(Z)_{ij}\neq\operatorname{\mathit{0}}\\ (<*)&\textbf{otherwise.}\end{cases}$
where $f^{\psi(z)}_{ij}$ are the structural equations of the edge $e_{ij}$ that are active under the MCS $T=\psi(z)$ and (in a slight imprecision in the actual definition) (<)absent(<*)( < ∗ ) is the function that carries on the previous function of the concatenation and discards $X_{j}$. While the ‘no-edge’ type $\operatorname{\mathit{0}}$ could be handled like any other type, we have listed it for clarity such that individual variables $\operatorname{\mathbf{X}}_{i}$ become contextually independent whenever $\psi(Z)_{ij}=\operatorname{\mathit{0}}$.

### v2

@@ 1 Introduction
Structural Causal Models (SCM) have become the de facto formalism in causality by representing causal relations as structural equation models (Pearl (2009); Spirtes et al. (2000)). While SCM are most useful for representing the intricate mechanistic details of systems, it can be challenging to derive the general qualitative behavior that emerges from the interplay of individual equations. Talking about the general type of relation emitted by particular mechanisms generalizes above the narrow computational view and has the ability to inspect causal systems from a more general perspective.
In addition to that, causal graphs can be subject to change whenever novel mechanisms emerge or vanish within a system. Prominently, agents can ‘break’ the natural unfolding of systems dynamics by forecasting system behavior and preemptively intervening in the course of events. As inherent parts of the environment, agents commonly establish or suppress the emergence of causal connections (Zhang & Bareinboim, 2017; Lee & Bareinboim, 2018; Dasgupta et al., 2019).
Consider the scenario shown in Figure 1 (left), where an agent $A$ (with position $A_{X}$) follows an agent $B$ (with position $B_{X}$) according to its internal policy $A_{\operatorname{\bm{\pi}}}$. We are interested in answering the question ‘What is the cause of the current position of agent A?’. In general, the observed system can be formalized as follows: $A_{X}:=f_{A}(B_{X},U_{A})$ and $B_{X}:=f_{B}(U_{B})$. Note that, from a classical causal perspective, we observe $B\rightarrow A$, since $A$ self-conditions itself to follow $B$, by instantiating the equation $f_{A}$ via its policy and thus becomes dependent on $B$. Classical causal considerations, which only consider how relations between variables are constructed, cannot give the correct answer. Only when we take a meta-causal stance and think about how equation $f_{A}$ came to be –how meta-causal states induce qualitative changes in behavior–, we can give a sufficient answer to this question.
@@ 2 Background
Causal Models. A common formalization of causality is provided via Structural Causal Models (SCM; (Spirtes et al., 2000; Pearl, 2009)). An SCM is defined as a tuple $\operatorname{\mathcal{M}}:=(\operatorname{\mathbf{U}},\operatorname{\mathbf{V}},\operatorname{\mathbf{F}},P_{\operatorname{\mathbf{U}}})$, where $\operatorname{\mathbf{U}}$ is the set of exogenous variables, $\operatorname{\mathbf{V}}$ is the set of endogenous variables, $\operatorname{\mathbf{F}}$ is the set of structural equations determining endogenous variables, and $P_{\operatorname{\mathbf{U}}}$ is the distribution over exogenous variables $\operatorname{\mathbf{U}}$. Every endogenous variable $V_{i}\in\operatorname{\mathbf{V}}$ is determined via a structural equation $v_{i}:=f_{i}(\text{pa}(v_{i}))$ that takes in a set of parent values $\text{pa}(v_{i})$, consisting of endogenous and exogenous variables, and outputs the value of $v_{i}$. The set of all variables is denoted by $\operatorname{\mathbf{X}}=\operatorname{\mathbf{U}}\cup\operatorname{\mathbf{V}}$ with values $\mathbf{x}\in\operatorname{\bm{\mathcal{X}}}$ and $N=|\operatorname{\mathbf{X}}|$. Every SCM $\operatorname{\mathcal{M}}$ entails a graph $\mathcal{G}^{\operatorname{\mathcal{M}}}$ that can be constructed by adding edges $(X_{i},X_{j})\in\operatorname{\mathbf{X}}\times\operatorname{\mathbf{X}}$ for each variable $X_{j}\in\operatorname{\mathbf{X}}$ and its parents $X_{i}\in\text{Pa}(X_{j})$. This can be expressed as an adjacency matrix $A\in\mathbb{B}^{N\times N}$ where $A_{ij}:=1$ if $(i,j)\in\mathcal{G}$ and $A_{ij}:=0$ otherwise.
@@
Mediation Processes. Causal effects are embedded in an environment that governs the dynamics and mediates between different causes and effects. To reason about the existence of causal relations, we need to consider this process-embedding environment. We define a mediation process $\operatorname{\mathcal{E}}=(\mathcal{S},\operatorname{\sigma})$, adapted from Markov Decision Processes (Bellman (1957)) for our setting. Here, $\mathcal{S}$ is the state space of the environment and $\operatorname{\sigma}:\mathcal{S}\rightarrow\mathcal{S}$ is the (possibly non-deterministic) transition function that takes the current state and outputs the next one. If we also have an initial state $\operatorname{s_{0}}\in\mathcal{S}$, we call $(\operatorname{\mathcal{E}},\operatorname{s_{0}})$ an initialized mediation process. As we are concerned with the general mediation process, we omit the common notion of a reward function $r$. Furthermore, we omit an explicit action space $\mathcal{A}$ and agents’ policy $\operatorname{\bm{\pi}}$ and model actions directly as part of the transition function $\operatorname{\sigma}$. In accordance to considerations of Cohen (2022), this eases treatment of environment processes and agent actions as now both are defined on the same domain. The emergence of SCM from mediation processes can be studied under a measure-theoretic theory, as considered in Park et al. (2023). Similarly, Janzing & Mejia (2022) discuss the role of elementary actions towards the constitution of causal variables.
@@
Typing. In the following section, we will make use of an identification function $\operatorname{\mathcal{I}}$ to determine the presence or absence of edges between any two variables. In particular, one can make use of different identification functions to identify different types of edges. Previous work on typing causality exists (Brouillard et al., 2022), but primarily considers the types of variables that are causally related, rather than the type of structural relation itself. Other works in the field of cognitive science consider the perception of different types of mechanistic relations (e.g., ‘causing’, ‘enabling’, ‘preventing’, ‘despite’) based on the role that different objects play in physical scenes (Chockler & Halpern, 2004; Wolff, 2007; Sloman et al., 2009; Walsh & Sloman, 2011; Gerstenberg, 2022). Since all objects are usually governed by the same physical equations, this assignment of types serves to provide post hoc explanations of a scene, rather than to identify inherent properties of its computational aspects. Gerstenberg (2024) provides an ‘intuitive psychology’ example that fits well with our scenario.
@@ 3 Meta-Causality
Similar to Pearl & Mackenzie (2018), we define Meta-Causality to be the [science of] change in qualitative cause-effect behavior. Usually, the mediating process may be too fine-grained to yield interpretable models. Therefore, we consider a set of variables of interest $\operatorname{\mathbf{X}}$ modeled by an SCM and introduced by an abstraction $\varphi:\operatorname{\mathcal{S}}\to\operatorname{\bm{\mathcal{X}}}$. (We provide a brief discussion on the emergence of SCM from mediating processes in Appendix A). Here, $\varphi$ could be defined as a summarization or causal abstraction over the state space (Rubenstein et al., 2017; Beckers & Halpern, 2019; Anand et al., 2022; Wahl et al., 2023; Kekić et al., 2023; Willig et al., 2023). In order to identify the type of causal relations from a mediating process, we need to be able to decide on what constitutes a type.
@@
For a given mediation process $\operatorname{\mathcal{E}}=(\operatorname{\mathcal{S}},\operatorname{\sigma})$ a meta-causal frame is a tuple $\operatorname{\mathcal{F}}=(\operatorname{\mathcal{E}},\operatorname{\mathbf{X}},(\operatorname{\tau}_{ij}),\operatorname{\mathcal{I}})$ with:
  • type-encoders $\operatorname{\tau}_{ij}:\operatorname{\mathcal{X}}_{i}\times\operatorname{\bm{\mathcal{X}}}^{\operatorname{\mathcal{S}}}\to\operatorname{\mathcal{T}}$ that assign a type $t\in\operatorname{\mathcal{T}}$ to the functional dependence of $X_{j}$ on $X_{i}$, induced by the underlying process $(\operatorname{\mathcal{S}},\operatorname{\sigma})$. This relation is between $\operatorname{\mathcal{X}}_{i}$ (values of $X_{i}$) and the abstraction of the transition function $\varphi\circ\operatorname{\sigma}\in\operatorname{\bm{\mathcal{X}}}^{\operatorname{\mathcal{S}}}=\{\psi:\operatorname{\mathcal{S}}\to\operatorname{\bm{\mathcal{X}}}\}$.
  • an identification function $\operatorname{\mathcal{I}}:\operatorname{\mathcal{S}}\times\operatorname{\mathbf{X}}\times\operatorname{\mathbf{X}}\rightarrow\operatorname{\mathcal{T}}$ with $\operatorname{\mathcal{I}}(s,X_{i},X_{j})\mapsto t:=\operatorname{\tau}_{ij}(\varphi(s),\varphi\circ\operatorname{\sigma})$ that assigns a type to every pair of causal variables for any state of the environment.
Types generalize the role of edges in causal graphs, while type encoders $\operatorname{\tau}_{ij}$ determine the particular type of edges from properties of the underlying functional relations $\varphi\circ\operatorname{\sigma}$. In most classical scenarios, the co-domain $\operatorname{\mathcal{T}}$ of the type encoder $\operatorname{\tau}_{ij}$ is chosen to be Boolean, representing the existence or absence of edges. In other cases, different values $t\in\mathcal{T}$ can be understood as particular types of edges, like positive, negative, or the absence of influence. This will help us to distinguish meta-causal states that share the same graph adjacency. The only requirement for $\operatorname{\mathcal{T}}$ is that it must contain a special value $\operatorname{\mathit{0}}$, which indicates the total absence of an edge. Meta-causal states now generalize the idea of binary adjacency matrices. We, intentionally, do not restrict the identification function in any way. However, particular choices, e.g. functions identifying only direct causal effects, are more likely to result in classical SCM. We provide a short discussion on this in Appendix B.
@@
In a meta-causal frame $\operatorname{\mathcal{F}}=(\operatorname{\mathcal{E}},\operatorname{\mathbf{X}},\operatorname{\tau}_{ij},\operatorname{\mathcal{I}})$, a meta-causal state is a matrix $T\in\operatorname{\mathcal{T}}^{N\times N}$. For a given environment state $s\in\operatorname{\mathcal{S}}$, the actual meta-causal state $T_{s}$ has the entries $T_{s,ij}:=\operatorname{\mathcal{I}}(s,X_{i},X_{j})=\operatorname{\tau}_{ij}(\varphi(s),\varphi\circ\operatorname{\sigma})$.
@@
Usually, we have the objective to learn the transition function $\delta$ for an unknown state transition function $\operatorname{\sigma}$. The state transition function $\delta$ can be approximated as $\delta(T,s):=\operatorname{\mathcal{I}}(\operatorname{\sigma}(T,s),X_{i},X_{j})=\operatorname{\tau}_{ij}(\varphi(s),\varphi\circ\operatorname{\sigma}(T,s))$.
@@
Inferring the Meta-Causal States. Even if the state transition function is known, it may be unclear from a single observation which exact meta-causal state led to the generation of a particular observation $S$. This is especially the case when two different meta-causal states can fit similar environmental dynamics. Even in the presence of latent factors (e.g., an agent’s internal policy), the current dynamics of a system (e.g., induced by the agent’s current policy) can sometimes be inferred from a series of observed environment states. This requires knowledge of the meta-causal dynamics, and is subject to the condition that sequences of observed states uniquely characterize the meta-causal state. Since MCM are defined as finite state machines, the exact condition for identifying the meta-causal state is that observed sequences are homing sequences (Moore et al., 1956). Note that the following example of a game of tag presented below exactly satisfies this condition, where the meta-causal states produce disjoint sets of environment states (‘agent A faces agent B’, or ‘agent B faces agent A’), and thus can be inferred from either a single observation, when movement directions are observed, or two observations, when they need to be inferred from the change in position of two successive observations.
@@
$(t_{B\rightarrow A}=\text{``A Chasing''})\iff(\dot{A}_{pos}\bm{\cdot}(B_{pos}-A_{pos})>0)$
@@
Causal and Anti-Causal Meta-Causal States. Assigning meta-causal states to particular system observations can be understood as labeling the individual observations. However, it is generally unclear whether the meta-causal state has an observational or a controlling character on the system under consideration. One could ask the question whether the system dynamics cause the meta-causal state, whether the meta-causal state causes the system dynamics, (or whether they are actually the same concept), similar to the well-known discussion “On Causal and Anticausal Learning” by Schölkopf et al. (2012), but from a meta-causal perspective.
@@
Role of Contextual Independencies. Changes in the causal graph are often attributed to changes in the environment modeled by some exogenous variables $Z$, as, for example, leveraged with the invariant causal prediction (Peters et al., 2016). In this case, we get MCM where the transition function only contains self-loops. This makes it clear that MCM are a more general tool in analyzing meta-causal structures. If we introduce a ‘no-edge’ type $t=\operatorname{\mathit{0}}$ for this scenario, a meta-causal model can describe the condition that $X_{i}$ and $X_{j}$ are contextually independent for some $Z=z$. For a suitable type-encoder with a surjective mapping $\psi:\mathcal{Z}\to\operatorname{\mathcal{T}}^{N\times N}$ and a given family of compatible (see Appendix C) decomposed structural equations $(f^{z}_{ij})_{z\in\mathcal{Z}}$ we have $f^{z}_{j}:=\bigcirc_{i\in 1..N}f_{ij}|Z$ and
$f_{ij}|Z:=\begin{cases}f^{z}_{ij}&\textbf{if~{}}\psi(z)_{ij}\neq\operatorname{\mathit{0}}\\ (<*)&\textbf{otherwise}\end{cases}$ (1)
where $f^{z}_{ij}$ are the structural equations of the edge $e_{ij}$ that are active under the environment $Z$ and (in a slight imprecision in the actual definition) (<)absent(<*)( < ∗ ) is the function that carries on the previous function of the composition and discards $X_{j}$. While the ‘no-edge’ type $\operatorname{\mathit{0}}$ could be handled like any other type, we have listed it for clarity such that individual variables $\operatorname{\mathbf{X}}_{i}$ become contextually independent whenever $f_{ij}|Z=\operatorname{\mathit{0}}$. We provide conditions for the reduction of MCM in Appendix D.